# DynamicScaler: Seamless and Scalable Video Generation for Panoramic Scenes

### v1

@@ Abstract
The increasing demand for immersive AR/VR applications and spatial intelligence has heightened the need to generate high-quality scene-level and 360° panoramic video. However, most video diffusion models are constrained by limited resolution and aspect ratio, which restricts their applicability to scene-level dynamic content synthesis. In this work, we propose the DynamicScaler, addressing these challenges by enabling spatially scalable and panoramic dynamic scene synthesis that preserves coherence across panoramic scenes of arbitrary size. Specifically, we introduce a Offset Shifting Denoiser, facilitating efficient, synchronous, and coherent denoising panoramic dynamic scenes via a diffusion model with fixed resolution through a seamless rotating Window, which ensures seamless boundary transitions and consistency across the entire panoramic space, accommodating varying resolutions and aspect ratios. Additionally, we employ a Global Motion Guidance mechanism to ensure both local detail fidelity and global motion continuity. Extensive experiments demonstrate our method achieves superior content and motion quality in panoramic scene-level video generation, offering a training-free, efficient, and scalable solution for immersive dynamic scene creation with constant VRAM consumption regardless of the output video resolution. Our project page is available at https://dynamic-scaler.pages.dev/
11footnotetext: Equal contribution.22footnotetext: Corresponding authors.
{strip}[Uncaptioned image]
Figure 1: We introduce DynamicScaler, a framework for generating dynamic panoramas conditioned on both images and text, or text alone. DynamicScaler enables the creation of arbitrary rectangular panoramas as well as 360° panoramic views, offering immersive visual experiences for AR/VR applications and displays of any size. (Please refer to our project page https://dynamic-scaler.pages.dev/ for better visualization.)
@@ 1 Introduction
Recent methods for image generation face two key challenges: generating high-resolution or wide aspect ratio images, and maintaining motion consistency and memory efficiency in dynamic scene generation, such as video synthesis. Extending image generation to higher resolutions or wider aspect ratios is computationally intensive, requiring significant memory and large-scale training datasets. For example, approaches like SDXL [18] support larger range of aspect ratios but still faces scalability issues, especially when it comes to ultra wide aspect ratio and higher resolution, which limits this methods in consistency, memory usage and inference speed. Other methods, such as those that stitch patches from pre-trained diffusion models, work well for generating panoramic or landscape images with repetitive patterns [13, 2], but this approach is less effective for more complex or varied scenes. Additionally, although there have been advances in spatially scalable diffusion models for static image generation, these models are often limited to square images [6, 8, 9, 30], restricting their ability to handle broader aspect ratios. In contrast, the challenge of dynamic scene generation requires not only spatial coherence across frames, but also global motion consistency, making it even more computationally demanding. Moreover, video generation models must be designed with memory efficiency in mind, as large-scale dynamic scene synthesis often strains memory, limiting real-time inference capabilities. While recent works have explored training-free approaches for expanding diffusion models to new domains [20, 12, 21], the scalability of these models to high-resolution video generation remains largely underexplored, requiring solutions that balance motion consistency and memory consumption effectively.
@@
  • Extensive experiments demonstrate that DynamicScaler outperforms existing methods in visual quality and motion consistency, generating continuous, loopable dynamic scenes suitable for immersive applications.
2 Related Work
@@
Spatial Scaling of Diffusion Models. Diffusion models have achieved remarkable success in generating high-quality images, with recent efforts focusing on scaling to diverse resolutions and aspect ratios. Existing approaches often rely on retraining or fine-tuning large models [31, 25], incurring significant computational costs and requiring extensive datasets. Other methods generate panoramic images by stitching patches from pretrained models, which work well for repetitive patterns like landscapes [13, 2, 7, 19]. Additionally, techniques such as AnyLens [20], MagicScroll [21], and AutoDiffusion [16] extend pretrained models across various domains without retraining. However, most of these methods are limited to square aspect ratios or fixed resolutions [6, 8, 9, 30] and cannot directly generate panoramic scenes.
@@
Panoramic Scene Synthesis. Recent advancements in scene-level generation focus on synthesizing large-scale 3D scenes from text, as seen in works like LucidDreamer, GALA3D, and Wonderworld [4, 32, 28]. These methods primarily generate static 3D scenes using Gaussian representations, limiting their capacity for dynamic content or flexible perspectives. Research on immersive generation has shifted toward 360° panoramas. OmniDreamer [1] employs cyclic inference for 360° image synthesis, while ImmenseGAN [5] leverages fine-tuning for better control. Diffusion-based methods, such as DiffCollage [29] and PanoDiff [22], have shown promise for static panoramas, yet fail to address dynamic video generation. While DreamScene360 [15] integrates Gaussian splatting for text-to-panorama synthesis, its reliance on static priors restricts dynamic scene applications.
@@
Dynamic Scene Generation. Dynamic scene generation for panoramic videos introduces challenges in maintaining motion coherence, temporal consistency, and visual quality over extended durations. 360DVD [23] adapts video diffusion models for panoramic data but faces limitations in generalization, interpolation accuracy, and style diversity when combining multiple LoRA models. Similarly, methods like 4K4DGen [17] and Vividdream [14] use overlapping regions for scene animation but suffer from noisy artifacts, fixed-window denoising, and restricted motion ranges. Moreover, 4K4DGen depends on inefficient optimization pipelines and is limited to image-to-video transformations. In contrast, our method overcomes these limitations with a shifting-based denoising framework that ensures spatial and temporal coherence. Our approach enables the generation of seamless, long-duration, and loopable panoramic videos with high-quality dynamic motion.
@@ 3 Method
Figure 3: The purposed Offset Shifting Window mechanism, which involves shifting denoising windows both vertically and horizontally between denoise steps to denoise the whole panorama video latent with arbitrary aspect ratio an resolution. Generally, all denosing windows are shifted vertically and horizontally every step, with different ways to handle the windows that cross the spatial boundary.
Figure 4: Results showcasing the scalability of our model across various resolutions, including 360° and rectangular panorama settings as well as both text conditioned and image conditioned generation. For more results please refer to our project page.
We present DynamicScaler, a scalable and tuning-free framework for panoramic dynamic scene synthesis that achieves seamless spatial and temporal coherence in dynamic environments. To extend video diffusion from fixed resolutions to expansive panoramas, our approach combines multiple diffusion model processes conditioned on noisy latents from previous timesteps, allowing denoising to adapt across spatial dimensions by adjusting the shifting path within panoramic space, as illustrated in Figure 2. To further enhance structural awareness and ensure global coherence in large-scale motion, we introduce Global Motion Guidance, a structured motion prior tailored for panoramic video generation. This guidance, integrated with Spatial Shifting Denoising, preserves both broad motion structures and intricate details, yielding cohesive and richly detailed scenes. Additionally, our framework provides methods for continuous, loopable motion, enabling seamless transitions and frame-to-frame consistency for extended, loopable video sequences.
@@ 3.1 Offset Shifting Denoising
We propose a dynamic scene synthesis approach integrating multiple reference diffusion processes within a unified framework. Existing methods such as SyncDiffusion [13] and MultiDiffusion [2] focus on image generation with patch-based techniques limited by local overlaps. However, these models suffer from restricted influence beyond adjacent areas, leading to blurred details and increased inference time due to the high overlap requirements.
The proposed DynamicScaler utilizes a shifting denoising strategy that dynamically adjusts the denoising process across the entire scene rather than being confined to fixed regions to address these limitations. This enables the synthesis of dynamic scenes, including 360° panoramic spaces, where the shifting denoising path plays a crucial role in montage-style and arbitrary-sized visual generation. By conditioning motion transfer on the interaction between noisy latents from different diffusion models, DynamicScaler allows for information sharing across regions, achieving seamless scene synthesis. By combining the denoising outputs from multiple models where each contributing unique spatial details, we ensure global consistency and significantly enhance visual fidelity. This approach improves the quality of dynamic scenes and scales flexibly to various scene dimensions and orientations.
In each steps, denoising windows are shifted both vertically and horizontally as illustrated in Figure 3. Specifically, in vertical shifting, the windows with in the boundaries are shifted vertically by an offset every steps, and the region that is not covered by shifting windows are handled by padding windows at the top and bottom. Horizontally, windows are also shifted by offset while the whole video latent would be regarded as a ring, that is, the left and right boundaries are connected so windows can cross through the boundary, which is implemented by filling the ’out of boundary’ regions by the corresponding region from the other side. This mechanism not only allows seamless transition across the whole generated video in arbitrarily aspect ratio but also enables motion and content continuity between the left and right boundaries, as shown in fig-5, producing immersive panoramic videos. The Offset Shifting Denoising process can be formulated as:
$Z_{t}=Con(\Phi_{\theta}(t,c,S(Z_{t-1},i,j,t)),t)$ (1)
where $S(X,i,j,t)$ indicates dividing the panorama video latent $X$ into $n_{w}\times n_{h}$ windows with offset at time $t$ and select the one in $i-th$ column, $j-th$ row. $Con(\cdot)$ means merging latent from divided windows back into $X$. $\Phi_{\theta}(t,c,x)$ represents a Denoising UNet, for T2V, $c=\{c_{text}\}$; and for I2V, $c=\{c_{text},S(c_{img},i,j,t)\}$.
@@ 3.2 Global Motion Guidance
Shifting denoising along horizontal and vertical directions enables scalable panorama generation and seamless motion. However, complex motion patterns, such as cascading waterfalls, require coordination across the entire panoramic field. To address this, we decompose the generation process into global layout and local content stages hierarchically: global layout stage captures high-level motion structures in lower resolution, while local content stage refines fine details and support higher resolution while maintain consistency. We apply causal global motion guidance to structure overall scene dynamics, followed by progressive upscaling to enhance resolution. Starting with a low-resolution generation, the process iteratively scales up by a factor of $K$, refining details through an “initialize-denoise” phase and subsequent “upsample-diffuse-denoise” loops. The upscaling process is formulated as:
$p_{\theta}(\mathbf{z}_{0}^{S}|\mathbf{z}_{T}^{1})=p_{\theta}(\mathbf{z}_{0}^{1}|\mathbf{z}_{T}^{1})\prod_{s=2}^{S}\left(q(\mathbf{z^{\prime}}_{T}^{s}|\mathbf{z^{\prime}}_{0}^{s})p_{\theta}(\mathbf{z}_{0}^{s}|\mathbf{z^{\prime}}_{T}^{s})\right),$ (2)
where $\mathbf{z}_{0}^{\prime s}=\text{inter}(\mathbf{z}_{0}^{s-1})$, with $\text{inter}(\cdot)$ denoting an interpolation algorithm like bicubic. This iterative process compensates for interpolation artifacts and refines details, ensuring coherent, high-quality panorama generation.
@@
Our framework enables the generation of 360° FoV panoramas, conditioned on text or images, using a sliding window denoising mechanism. This ensures that the resulting panorama covers the full 360° field of view, making it suitable for immersive applications. For rectangular panoramas, the sliding window shifts across both rows and columns to improve scene coherence. For 360° FoV panoramas in equirectangular projections (ERPs), the image is first reprojected into spherical space, where the sliding window mechanism is applied to ensure full coverage of the entire sphere. However, interpolation in equirectangular projections can introduce distortion, particularly in large polar angles. To mitigate this, we apply equirectangular interpolation during the early denoising steps to capture the broad structure of the scene, and later shift in spherical space to refine finer details and preserve motion fidelity. This two-stage process ensures high-quality 360° FoV panorama generation with minimal distortion.
360° FoV Panoramic Representation.
Panoramic images in equirectangular format are represented as an $H\times W\times C$ matrix. To minimize distortion, we use a spherical representation where pixel values are defined on the unit sphere $S^{2}$. In spherical coordinates, a point on the unit sphere is represented by the directional vector $d=(x,y,z)$, where $x=\sin(\theta)\cos(\phi)$, $y=\sin(\theta)\sin(\phi)$, and $z=\cos(\theta)$, with $\theta$ being the polar angle (latitude) and $\phi$ being the azimuthal angle (longitude). The pixel value at direction $d$ is then mapped as follows:
$S_{I}(d)=E_{I}\left(\frac{1}{\pi}\arccos(y),\frac{1}{2\pi}\arcsin(z)\right)$ (3)
where $S_{I}(d)$ denotes the pixel value at direction $d=(x,y,z)$. Here, the function $E_{I}$ maps the spherical coordinates to the corresponding pixel values in the equirectangular image.
Offset Shifting in Panoramic Space.
Most video diffusion models are trained by datasets in perspective views, thus can not be directly applied to generate spherical 360 degree panoramic video, which is usually represented in equirectangular projection.
Figure 5: The purposed Panoramic Projecting Denoise, where spherical panorama videos (represented as equirectangular projections) are denoised by projecting them into perspective view windows, followed by re-projection back to the equirectangular format, which allowed denoising 360 degree panoramic video with diffusion models trained in regular (perspective) views dataset.
To address this, We apply Panoramic Projecting Denoise to spherical panorama videos, as illustrated in Figure 5 The panorama is represented in equirectangular projection, then projected onto perspective windows, denoised, and re-projected back at every denoising steps. The window angles are adjusted through yaw rotations at each step to ensure spatial coherence across the entire panorama, which can be formulated as:
$Z_{t}=\tilde{V}_{c}\left(\Phi_{\theta}\left(t,c,V(Z_{t-1},a_{i,j,t},f)\right),t\right)$ (4)
where $V(X,a,f)$ projects the panorama latent $Z$ in ERP format at view angle $a$ and fov $f$ into perspective views, similar to regular OSD, the view angles $a$ have an offset at each timestep $t$, creating perspective view windows that can be denoised using normal diffusion models $\Phi_{\theta}$. $\tilde{V}_{c}$ projects all the view windows back into ERP, updating the whole panorama latent. Similar with regular perspective Offset Shifting Denosing, we can also shift the view windows each steps, with an offset applied in pitch angles to denoise the whole video seamlessly, allowing us to generate spherical 360 degree panoramic video by regular diffusion models without extra training.
@@
In cases that windows have overlap with each others, which is often the case in high polar angle regions, parts of the latent space corresponding to subsequent windows may have already been denoised by previous windows, creating inconsistent noise levels. To address this, the system maintains a denoising mask $M_{d}$ at each timestep $t$ to track denoised regions. Before processing each window, noise is rebalanced as:
$Z^{\prime}_{t}=\begin{cases}\sqrt{\alpha_{t}}Z_{t}+\sqrt{1-\alpha_{t}}\epsilon_{t}&\text{if }M_{d}(x,y)=1\\ Z_{t}&\text{if }M_{d}(x,y)=0\end{cases}$ (5)
@@
ensuring consistent noise levels across all windows before passing into diffusion model. More details can seen from the appendix.

### v2

@@ Abstract
The increasing demand for immersive AR/VR applications and spatial intelligence has heightened the need to generate high-quality scene-level and 360° panoramic video. However, most video diffusion models are constrained by limited resolution and aspect ratio, which restricts their applicability to scene-level dynamic content synthesis. In this work, we propose DynamicScaler, addressing these challenges by enabling spatially scalable and panoramic dynamic scene synthesis that preserves coherence across panoramic scenes of arbitrary size. Specifically, we introduce a Offset Shifting Denoiser, facilitating efficient, synchronous, and coherent denoising panoramic dynamic scenes via a diffusion model with fixed resolution through a seamless rotating Window, which ensures seamless boundary transitions and consistency across the entire panoramic space, accommodating varying resolutions and aspect ratios. Additionally, we employ a Global Motion Guidance mechanism to ensure both local detail fidelity and global motion continuity. Extensive experiments demonstrate our method achieves superior content and motion quality in panoramic scene-level video generation, offering a training-free, efficient, and scalable solution for immersive dynamic scene creation with constant VRAM consumption regardless of the output video resolution. Project page is available at https://dynamic-scaler.pages.dev/new .
Figure 1: We introduce DynamicScaler, a framework for generating dynamic panoramas conditioned on both images and text, or text alone. DynamicScaler enables the creation of regular panoramas with arbitary aspect ratio as well as 360° panoramic views, offering immersive visual experiences for AR/VR applications and displays of arbitrary aspect ratio and resolution.
@@ 1 Introduction
Recent methods for image generation face two key challenges: generating high-resolution or wide aspect ratio images, and maintaining motion consistency and memory efficiency in dynamic scene generation, such as video synthesis. Extending image generation to higher resolutions or wider aspect ratios is computationally intensive, requiring significant memory and large-scale training datasets.
Models like  [20] support larger range of aspect ratios but still faces scalability issues, especially when it comes to ultra wide aspect ratio and higher resolution, which limits this methods in consistency, memory usage and inference speed. Other methods, such as those that stitch patches from pre-trained diffusion models, work well for generating panoramic or landscape images with repetitive patterns [14, 2], but this approach is less effective for more complex or varied scenes. Additionally, although there have been advances in spatially scalable diffusion models for static image generation, these models are often limited to square images [7, 9, 10, 31], restricting their ability to handle broader aspect ratios. In contrast, the challenge of dynamic scene generation requires not only spatial coherence across frames, but also global motion consistency, making it even more computationally demanding. Moreover, video generation models must be designed with memory efficiency in mind, as large-scale dynamic scene synthesis often strains memory, limiting real-time inference capabilities. While recent works have explored training-free approaches for expanding diffusion models to new domains [22, 13, 23], the scalability of these models to high-resolution video generation remains largely underexplored, requiring solutions that balance motion consistency and memory consumption effectively.
@@
  • Extensive experiments demonstrate that DynamicScaler outperforms existing methods in visual quality and motion consistency, generating continuous, longer duration an loopable dynamic scenes suitable for immersive applications.
Figure 3: The purposed Offset Shifting Window mechanism, which involves shifting denoising windows both vertically and horizontally between denoise steps to denoise the whole panorama video latent with arbitrary aspect ratio an resolution. The denosing windows are shifted vertically and horizontally every step, creating ”overlap” regions between steps which mitigate the artifacts and synchronize the whole denoising process across the panorama. This results in seamless and consistent panoramic video generation with high resolution and aspect ratio.
2 Related Works
2.1 Spatial Scaling of Diffusion Models
@@
Diffusion models have achieved remarkable success in generating high-quality images, with recent efforts focusing on scaling to diverse resolutions and aspect ratios. Existing approaches often rely on retraining or fine-tuning large models [32, 27], incurring significant computational costs and requiring extensive datasets. Other methods generate panoramic images by stitching patches from pretrained models, which work well for repetitive patterns like landscapes [14, 2, 8, 21]. Additionally, techniques such as AnyLens [22], MagicScroll [23], and AutoDiffusion [17] extend pretrained models across various domains without retraining. However, most of these methods are limited to square aspect ratios or fixed resolutions [7, 9, 10, 31] and cannot directly generate panoramic scenes with large aspect ratio.
2.2 Panoramic Scene Synthesis.
@@
Recent advancements in scene-level generation focus on synthesizing large-scale 3D scenes from text, as seen in works like LucidDreamer, GALA3D, and Wonderworld [5, 33, 29]. These methods primarily generate static 3D scenes using Gaussian representations, limiting their capacity for dynamic content or flexible perspectives. Research on immersive generation has shifted toward 360° panoramas. OmniDreamer [1] employs cyclic inference for 360° image synthesis, while ImmenseGAN [6] leverages fine-tuning for better control. Diffusion-based methods, such as DiffCollage [30] and PanoDiff [24], have shown promise for static panoramas, yet fail to address dynamic video generation. While DreamScene360 [16] integrates Gaussian splatting for text-to-panorama synthesis, its reliance on static priors restricts dynamic scene applications.
2.3 Dynamic Scene Generation.
@@
Dynamic scene generation for panoramic videos introduces challenges in maintaining motion coherence, temporal consistency, and visual quality over extended durations. 360DVD [25] adapts video diffusion models for panoramic data but faces limitations in generalization, interpolation accuracy, and style diversity when combining multiple LoRA models. Similarly, methods like 4K4DGen [18] and Vividdream [15] use overlapping regions for scene animation but suffer from noisy artifacts, fixed-window denoising, and restricted motion ranges. Moreover, 4K4DGen depends on optimization process to maintain coherence and is limited to image-to-video transformations. In contrast, our method overcomes these limitations with a shifting-based denoising framework that ensures spatial and temporal coherence. Our approach enables the generation of seamless, long-duration, and loopable panoramic videos with high-quality dynamic motion.
@@ 3 Method
We present DynamicScaler, a scalable and tuning-free framework for panoramic dynamic scene synthesis that achieves seamless spatial and temporal coherence in panoramic sythesis, as illustrated in Figure 2. To extend video diffusion from fixed resolutions to expansive panoramas, our approach involves an Offset Shifting Denoising mechanism. To further enhance structural awareness and ensure global coherence in large-scale motion, we introduce Global Motion Guidance, a structured motion prior tailored for panoramic video generation. Additionally, our framework provides methods for continuous, loopable motion, enabling seamless transitions and frame-to-frame consistency for extended, loopable video sequences.
@@ 3.1 Offset Shifting Denoising
To generate a $W_{p}\times H_{p}$ panorama $Z$ with a regular diffusion model $\theta$ trained in resolution $W_{\theta}\times H_{\theta}(W_{\theta}<W_{p},H_{\theta}<H_{p})$ without extra finetuning, $Z$ should be divided into $n_{W}\times n_{H}$ windows to fit in the applicable size and denoised. Existing methods in panorama image synthesis like MultiDiffusion [2] and SyncDiffusion [14] utilize more windows $(n_{H}>H_{p}/H_{\theta},n_{W}>W_{p}/W_{\theta})$ that overlap with each others to mitigate the seam and distortion caused by divided denoising windows, resulting in much higher computation overhead, while  [8] randomly shift the windows to reduce the overlap needed, those methods mainly focus on regular perspective panorama image. However, in high resolution 360 degree panoramic video synthesis require extending both vertically and horizontally, contents also differs between each windows, hindering one from directly applying those methods. To tackle this challenge, we introduce Offset Shifting Denoising (OSD) that shift denoising windows by an offset, both vertically and horizontally, across the entire panorama in each steps as illustrated in Figure 3. This offset creates ”overlap” regions among windows between denoise steps, synchronizing content and motions across windows, and the trend of discontinuity at the edge of windows in one certain denoise step are also mitigated at the next step because the edges are also shifted along with the window.
Specifically, in the vertical direction, the windows with in the upper and lower boundaries are shifted vertically by an offset every steps, and the region that is not covered by those shifting windows are handled by the padding windows at the top and bottom of the panorama. Horizontally, windows are also shifted by offset while the whole latent would be regarded as a ring, that is, the left and right boundaries are ’connected’ so windows can cross through those boundary, which is implemented by filling the ’out of boundary’ regions by the corresponding region from the other side. The Offset Shifting Denoising process can be formulated as:
$Z_{t}=Con|_{1:n_{W},1:n_{H}}(\Phi_{\theta}(t,c,Split(Z_{t-1},i,j,t,n_{W},n_{H})))$ (1)
where $Split(X,i,j,t,n_{W},n_{H})$ indicates dividing the panorama video latent $X$ into $n_{W}\times n_{H}$ windows with offset at time $t$ and select the one in ithi-th column, jthj-th row; $Con(\cdot)$ means collecting latent from divided windows and concat them to form the entire panorama latent; $\Phi_{\theta}(t,c,x)$ represents the diffusion process, for T2V, $c=\{c_{text}\}$; and for I2V, $c=\{c_{text},Split(c_{img},i,j,t)\}$.
OSD not only allows seamless transition across the whole generated video in arbitrarily aspect ratio but also enables motion and content continuity between the left and right boundaries, as shown in Figure 7, producing immersive panoramic videos.
@@ 3.2 Global Motion Guidance
Shifting denoising along horizontal and vertical directions enables scalable panorama generation and seamless motion. However, complex motion patterns, such as cascading waterfalls, require coordination across the entire panoramic field. At early denoising steps, the overall layout are constructed  [19], while the intrinsic synchronization offered by OSD has not accumulate enough influence in this period, different part of the panorama may tend to separate motion pattern, resulting in less consist overall motion, especially in high resolution panoramic scene generation. Besides applying more windows to create explicit spatial overlap in the first $K$ steps, we introduce the Global Motion Guidance (GMG), which decompose the generation process into global layout and local content stages hierarchically: the first stage synthesis a video in lower resolution, capturing high-level motion structures. Then in the upscaling stage those low-resolution result are upsampled using interpolation algorithm ${inter}(\cdot)$ like bicubic and re-noised ${noise}(\cdot)$, serving as an initialization that guide content layout and motion in the following high-resolution generation, which refines local details while preserve consistency. This process is formulated as:
${Z_{HR^{0}}}={\Phi_{\theta}}^{OSD}({noise}({inter}({\Phi_{\theta}}^{OSD}(Z_{LR^{T}}))))$ (2)
Integrated with Offset Shifting Denoise, Global Motion Guidance preserves both broad motion structures and intricate details, yielding cohesive and richly detailed high resolution panoramic scenes.
Figure 4: The purposed Panoramic Projecting Denoise, where spherical panorama videos latents (represented as equirectangular projections) are projected into perspective view port windows and denoised, followed by re-projection back to the equirectangular panorama, as shown in the upper part of the figure. Those view port windows are also shifted with an offset applied in their view angles at each steps, as shown in the lower part of the figure. For legibility, only a proportion of view port regions are shown in the figure.
@@
360° FoV panoramas are usually represented by its equirectangular projection (ERP), which project the spherical panorama into an $H\times W\times C$ matrix, with $W/H=2$. Common diffusion models trained by regular (perspective) video datasets face challenge in generating 360° FoV panorama videos due to the deformation in equirectangular projection. Besides training adapters with extra 360° FoV datasets  [25], projecting portion of the equirectangular panorama back into mulitple perspective view ports allows using regular diffusion models to denoise.  [18] relys on an optimization procedure to synchronize the diffusion process among each view ports, while it is constrained in Image to Video Generation and suffer from insufficient motion. We instead intergrate our offset shifting denoising mechanism with equirectangular projecting, ensuring high-quality 360° FoV panorama generation with minimal distortion.
3.3.1 Spherical Projection.
Given a point $(x,y,z)$ on the unit sphere, the corresponding longitude $\alpha$ and latitude $\beta$ can be computed by:
$\alpha=\arctan 2(y,x),\;\beta=\arcsin(z).$ (3)
The equirectangular projection maps these spherical coordinates $(\alpha,\beta)$ to a rectangular plane $(u,v)$ with $u=\alpha,v=\beta$, where the rectangular plane has a width of $2\pi$ and a height of $\pi$, maintaining a $1:2$ aspect ratio. For the inverse mapping, given a point $(u,v)$ on the rectangular plane, the corresponding 3D coordinates $(x,y,z)$ on the unit sphere can be computed by:
$x=\cos(\beta)\cos(\alpha),\;y=\cos(\beta)\sin(\alpha),\;z=\sin(\beta).$ (4)
3.3.2 Offset Shifting in Panoramic Space.
Perspective view ports at certain view angle $a$ and field of view $f$ can be obtained by the projecting function $Proj(\cdot)$ from the equirectangular projection $Z^{p}$ of a 360 degree panorama, dividing the panorama into $n_{\alpha}\times n_{\beta}$ view port windows that can be denoised using normal diffusion model $\theta$. Similar to the Offset Shifting in perspective panorama, we can also shift the view port windows at each steps with an offset applied in view angle $a$, as illustrated in Figure 4, which can be formulated as:
$Z_{t}=Con_{Proj}|_{1:n_{\alpha},1:n_{\beta}}(\Phi_{\theta}(t,c,Proj(Z_{t-1},f,a_{i,j,t})))$ (5)
where $Con_{Proj}(\cdot)$ means collecting latent from divided view port windows and project into the 360 degree panorama.
@@
In cases that windows have overlap with each others, which is often the case in high polar angle regions, parts of the latent space corresponding to subsequent windows may have already been denoised by previous windows, creating inconsistent noise levels. To address this, we maintains a mask $M_{d}$ at each timestep $t$ to track denoised regions. Before processing each window, noise is rebalanced as:
$Z^{\prime}_{t}[x,y]=\begin{cases}\sqrt{\alpha_{t}}Z_{t}[x,y]+\sqrt{1-\alpha_{t}}\epsilon_{t}&\text{if }M_{d}(x,y)=1\\ Z_{t}[x,y]&\text{if }M_{d}(x,y)=0\end{cases}$ (6)
@@
ensuring consistent noise levels across all windows before passing into diffusion model.
Figure 5: The Offset Shifting Denoising mechanism extended to temporal dimension. The upper part shows how the frame clip windows are shifted with an offset along the temporal dimension, with padding windows at the start and end of the video sequence. The lower part shows the loopable offset shifting denoising, with looping frames at the start and end of the frames sequence.
3.4 Long and Loopable Scene Video Generation